# Tracking Any Object Amodally

### v2

@@ 6 Discussion
Table 9: Number of MLP layers in Amodal Expander. Empirically, using a lightweight 2-layer MLP amodal expander proves sufficient to generate reasonable amodal detection results. All models were trained on TAO-Amodal training set for 20k iterations.
@@
In this appendix, we extend our discussion of the proposed dataset and method within the context of tracking any object with amodal perception. Specifically, we discuss details about the training and PasteNOcclude technique in Section 7, provide further empirical analysis in Section 8, and outline the annotation guidelines of our dataset in Section 9. We further provide comprehensive video demonstration of our dataset and qualitative results at https://tao-amodal.github.io.
7 Implementation details
Training Amodal Expander.
We trained amodal expander on TAO-Amodal training set for 20k iterations for all experiments unless specified. We used a 2-layer MLP as the architecture. The hidden dimension of MLP is selected as 256. ReLU [1] and dropout [48] with a probability of 0.2 were applied to each layer except the last one. Note that amodal expander is compatible with any existing detector or tracker. However, in our experiments, we implemented the expander in conjunction with GTR [64]. Additional architecture details of GTR align with the selection in [64]. We used 0.01 as the base learning rate and applied WarmupCosineLR [24] as the scheduler. The optimizer is selected as AdamW [40]. The batch size for training is 4. We adopted the training methodology outlined in [64], treating each image as an independent sequence. Data augmentation [63], which includes random cropping and resizing, was applied to each image to produce synthetic videos with a length of 8 frames. Beyond this, we further applied PasteNOcclude, introduced in Sec. 4.2 in the main paper, on top of the synthetic videos to automatically generate more occlusion scenarios. We provide the hyperparameter details of PasteNOcclude in the next section.
PasteNOcclude (PnO).
The visual example of PnO is illustrated in Fig. 6. To collect the segments as occluders from LVIS [21] and COCO [38], we utilize segmentation mask to mask out the background area and employ the bounding box to crop the object. We filter out those segments where the mask area is less than 70% of the bounding box area to ensure the object (occluder) is not occluded. In the training process, we view each image as a sequence and create an 8-frame sequence employing the data augmentation strategy outlined in [64] based on each single image. Subsequently, we randomly select 1 to 7 segments from the collection and place them at random locations. Additionally, we randomly adjust the height and width of the inserted segments within the range of $[12,192]$. To maintain smooth transitions between consecutive frames, we randomly determine the object’s location and size only in the first and last frames. The size and location in intermediate frames are obtained through interpolation.
8 More empirical analysis
We used the same evaluation metrics defined in Section 5.1 in the main paper. For easy reference, we summarize all the definitions in Table 10. We presented additional experiments involving state-of-the-art trackers in Section 8.1 and the amodal expander in Section 8.2.
8.1 Benchmarking off-the-shelf-trackers
Evaluation with higher IoU thresholds.
In Table 6, we evaluate the trackers with average precision (AP) averaged over 10 IoU thresholds from 0.5 to 0.95 at a step 0.05. The performance trend basically aligns with what we observed in Table 2 in the main paper. GTR [64] obtained strong performance in both detection and tracking. When evaluated with higher IoU thresholds, ViTDet [37] and SORT [4] demonstrate inferior detection performance compared to GTR, indicating a contrasting outcome compared to the results obtained at a 0.5 threshold. This shows the limitations of SORT [4] in accurately estimating bounding boxes.
Running trackers at higher fps.
We reported the performance of state-of-the-art trackers running at 5 fps in Tab. 7. We noticed that ViTDet [37] along with SORT [4] achieved the best performance among all the trackers. This aligns with our intuition as SORT estimates the location in the current frame based on prior-frame locations. This property benefits from running at higher fps. Nevertheless, achieving this performance improvement requires processing ViTDet on five times more frames than models operating at 1 fps, leading to increased computational demands.
8.2 Amodal expander experiments
Investigating key information for amodal box inference.
Table 8 reports different input choices to the amodal expander. Modal box (deltas) $\Delta b$, output by the regression head as shown in Fig. 2 in the main paper, are used to yield final modal box predictions when applied to region proposals and thus contain information about the exact location of modal box predictions. Proposal features encompass visual appearance information of the detected region proposals. Absence of visual cues significantly diminishes the performance of both detection and tracking under occlusion. Interestingly, the amodal expander, incorporating both modal delta and proposal features, yielded the most favorable outcomes, indicating that, beyond visual information, estimating modal box locations also contributes to effective amodal reasoning.
Number of MLP layers.
We tested with the depth of amodal expander architecture in Table 9. We observe a reverse-U pattern concerning the number of MLP layers, with two-layer MLPs demonstrating superior performance compared to other models. A one-layer MLP proves suboptimal in both detection and tracking. Notably, using a 1-layer MLP results in slightly inferior outcomes compared to fine-tuning the regression head, as indicated in Table 3 in the main paper. Our contention is that the regression head may derive benefits from pre-training on modal benchmarks.

### v3

@@ 6 Discussion
In this appendix, we extend our discussion of the proposed dataset and method within the context of tracking any object with amodal perception. Specifically, we discuss details about the training and PasteNOcclude technique in Section 7, provide further empirical analysis in Section 8, and outline the annotation guidelines of our dataset in Section 9. We further provide comprehensive video demonstration of our dataset and qualitative results at webpage/index.html.
7 Implementation details
7.0.1 Training Amodal Expander.
We trained amodal expander on TAO-Amodal training set for 20k iterations for all experiments unless specified. We used a 2-layer MLP as the architecture. The hidden dimension of MLP is 256. We apply ReLU [1] and dropout [51] with a probability of 0.2 to each layer except the last one. We implemented the expander in conjunction with GTR [71]. Architecture details of GTR align with the selection in the prior work [71]. We used 0.01 as the base learning rate and applied WarmupCosineLR [26] as the scheduler. The optimizer is AdamW [43]. The batch size for training is 4. We adopted the training methodology outlined in [71], treating each image as an independent sequence. We applied data augmentation [70], including random cropping and resizing, to each image to produce synthetic videos with a length of 8 frames. Beyond this, we further applied PasteNOcclude, introduced in Sec. 4.2 in the main paper, on top of the synthetic videos to automatically generate more occlusion scenarios. We provide the hyperparameter details of PasteNOcclude in the next section.
PasteNOcclude (PnO). We illustrated visual examples of PnO in Fig. 9. We mask the background area with the segmentation mask and collect the cropped object from LVIS [23] and COCO [41] to serve as occluders. We filter out segments where the mask area is less than 70% of the bounding box area to ensure that the occluder is not occluded. In the training process, we view each image as a sequence and create an 8-frame sequence employing the data augmentation strategy in GTR [71] based on each image. Subsequently, we randomly select 1 to 7 segments from the collection and place them at random locations. Further, we randomly adjust the height and width of the inserted segments within the range of $[12,192]$. We randomly determine the object’s location and size only in the first and last frames to ensure smooth transitions between consecutive frames. The size and location in intermediate frames are obtained through interpolation.
8 More empirical analysis
We used the evaluation metrics defined in Sec. 5.1 in the main draft. We summarize all the definitions in Table 13. We presented additional experiments involving state-of-the-art trackers in Section 13 and the amodal expander in Section 8.2.
8.1 Benchmarking off-the-shelf-trackers
8.1.1 Evaluation on TAO-Amodal validation set.
We report detection and tracking average precision (AP) numbers of SOTA off-the-shelf trackers on TAO-Amodal validation set running at 1fps with an IoU threshold 0.5 in Tab. 7. We also observed similar performance trends when running at 5fps with higher IoU thresholds, shown in Tabs. 9 and 8. Every off-the-shelf tracker was trained on either TAO [10] or LVIS [23], ensuring alignment of category vocabulary with our dataset as detailed in Sec. 5.1. We reproduced AOA [12] using their released implementation, with object detector trained on LVIS and tracking ReID head trained on TAO.
8.1.2 Using off-the-shelf modal trackers for amodal perception.
Table 7 reveals notable differences in detection AP between modal (Modal AP) and amodal annotations (AP), amounting to an 8.49% difference. Additionally, the amodal tracking AP experiences a substantial decline compared to modal tracking AP. These results highlight the difference between amodal and modal perception.
8.1.3 How well do standard trackers handle occlusion?
Existing off-the-shelf trackers exhibit reasonable performance in detecting non-occluded objects, with ViTDet achieving 39.71% AP[0.8,1] as revealed in Table 7. However, all trackers face challenges in handling heavily occluded, partially occluded (AP[0.1,0.8]) and out-of-frame (OoF) scenarios. We noticed that ViTDet operating at 5 fps benefits from the property of SORT to estimate the location in the current frame using past information in Tab. 9. Nevertheless, this improvement comes at the cost of processing ViT-Det on 5x more frames than models running at 1 fps. In contrast, amodal completion could be a promising way for efficiently handling occlusion.
@@ 8.1.3 How well do standard trackers handle occlusion?
Evaluation with higher IoU thresholds. In Table 8, we evaluate the trackers with average precision (AP) averaged over 10 IoU thresholds from 0.5 to 0.95 at a step 0.05. The performance trend basically aligns with what we observed in Table 7 in the main paper. GTR [71] obtained strong performance in both detection and tracking. When evaluated with higher IoU thresholds, ViTDet [40] and SORT [4] demonstrate inferior detection performance compared to GTR, indicating a contrasting outcome compared to the results obtained at a 0.5 threshold. This shows the limitations of SORT [4] in accurately estimating bounding boxes.
Running trackers at higher fps. We reported the performance of state-of-the-art trackers running at 5 fps in Tab. 9. We noticed that ViTDet [40] along with SORT [4] achieved the best performance among all the trackers. This aligns with our intuition as SORT estimates the location in the current frame based on prior-frame locations. This property benefits from running at higher fps, but it requires processing ViTDet on 5×more frames than models operating at 1 fps, heavily increasing computational demands.
8.2 Amodal expander experiments
8.2.1 Scaling up training data.
In Tab. 10, we scale up the training data to 4x by including test videos as train set and evaluate the amodal expander on the validation set. We note that simply increasing the size of the training data does not significantly improve the metrics compared to results shown in Tab. 2. This validates our design to propose TAO-Amodal as an evaluation benchmark.
8.2.2 Investigating key information for amodal box inference.
Table 11 reports different input choices to the amodal expander. Modal box (deltas) $\Delta b$, output by the regression head as shown in Fig. 5 in the main paper, are used to yield final modal box predictions when applied to region proposals and thus contain information about the exact location of modal box predictions. Proposal features includes visual appearance information of the detected region proposals. Absence of visual cues significantly diminishes the performance of both detection and tracking under occlusion. Interestingly, the amodal expander, incorporating both modal delta and proposal features, yielded the most favorable outcomes. This indicates that estimating modal box locations also contributes to effective amodal reasoning.
8.2.3 Number of MLP layers.
We tested with the depth of amodal expander architecture in Table 12. We observe a reverse-U pattern concerning the number of MLP layers, with two-layer MLPs demonstrating superior performance compared to other models. A one-layer MLP proves suboptimal in both detection and tracking. Notably, using a 1-layer MLP results in slightly inferior outcomes compared to fine-tuning the regression head, as indicated in Table 3 in the main paper. We argue that the regression head may derive benefits from pre-training on modal benchmarks.